esmirb:

The attendance at a seminar in year 1 was 500. In Year 2, the attendance changed by 100. What was the percent change in attendance?

kittybasil:

Original attendance: \(500\) New attendance: increases by 100, aka \(500+100=600\) Percent difference is \(\frac{\text{new}-\text{original}}{\text{original}}\), which means\[\frac{600-500}{500}=\frac{100}{500}=\frac{1}{5}\] \(1/5\) also translates to \(20/100\) or 20%. So your percent change is twenty percent.